Solve and check each of the following equations for x.
x/6-x/8=1
please help.
Thank you so much.

Answer by ikdeep(226)

x/6 - x/8 = 1
1st take the LCM of denominator i.e 6 & 8 which gives LCM = 24
on applying the LCM steps we get...
(4x -3x)/ 24 = 1
now on cross multiplying ..we get
4x - 3x = 24
x = 24

......to check you answer ,,,substitute the value of x = 24 in the given equation and we get
x/6 - x/8 = 1
24/6 - 24/8 = 1
4 - 3 = 1
1 = 1
LHS = RHS
Here checked or verified the answer

Please feel free to revert back for any further queries.